# reference.bib

### v1

@@ VII Simulation Studies
In this section, we conduct simulation studies to evaluate the performance of the proposed approaches. We start by describing the experiment setup in Sec. VII-A. Next, we conduct two sets of studies to evaluate the optimality and efficiency of the proposed approaches in Sec. VII-B and Sec. VII-C, respectively. Lastly, we investigate the impact of key parameters in Sec. VII-D.
@@ VII-A Experiment Setup
We evaluate the proposed approaches on different network graphs generated using the method introduced in [erdds1959random]. The network graphs are then transformed into tree topologies by using Dijkstra’s algorithm. In each network topology, we configure the computing capacities $f_{i}$ of the servers by randomly generating values from the range of $[1,10]$MHz. The values of the data transmission rates $R_{ij}$ are randomly generated from the range of $[10,100]$Mbps. Moreover, we set $\gamma_{i}=10^{-2}$, and $e_{i}=30$dBm for all $i\in\mathcal{N}$. The task size is set to $Y=1$ Gbits and $b=10^{6}$ cycles/Gbit. All approaches are implemented in Python and evaluated on an Alienware Aurora R15 Gaming Desktop with a 12 Gen Intel i9 CPU and 64G of memory.
@@ VII-B Optimality Analysis
In the second experiment, we randomly set the weights to $w_{1}=0.5$ and $w_{2}=0.0416$, so that both computation efficiency and energy consumption are considered in our approaches. Note that these weight values are also used in the following experiments. Fig. 5(b) shows the comparison results, demonstrating the promising performance of our approaches in balancing task completion time and energy consumption. In Fig. 6(a) and Fig. 6(b), we show the task completion time, i.e., $\max_{i\in\mathcal{N}}T_{i}^{total}$, and the maximum energy consumption by any server, i.e., $\max_{i\in\mathcal{N}}E_{i}^{total}$, respectively. The result indicates that our approach outperforms all the benchmarks in both task completion time and maximum energy consumption.
Figure 5: Total cost $J$ of different methods when (a) $w_{1}=1$, $w_{2}=0$; and (b) $w_{1}=0.5$, $w_{2}=0.0416$.
Figure 6: (a) Task completion time (b) Maximum energy consumption of different methods when $w_{1}=0.5$, $w_{2}=0.0416$.
@@ VII-C1 Small-Scale Networks
The base-10 logarithm of the execution time, i.e., $\log_{10}T_{\text{exe}}$, of each method is shown in Fig. 7(b). As expected, the optimal methods, CMO and PMO, are more time-consuming than the three heuristic methods. Moreover, PMO, being a parallelized version of CMO, significantly reduces execution time in Topologies 2-4 due to its parallelism. For Topology 1, since the root has a single subtree, PMO is equivalent to CMO. Among the heuristic methods, LP achieves the least execution time by pruning the most nodes and significantly reducing the search space. GA, on the other hand, is the least efficient and even underperforms PMO in Topologies 2 & 4. This suggests that for small networks, PMO can be directly applied. Furthermore, comparing NP and PMO, we can observe that NP does not improve efficiency in most scenarios, despite reducing the number of workers. This is because only one node is “pruned” in each topology, and the time saved by pruning is offset by the overhead generated by the pruning procedure.
@@ VII-D Parameter Impact Analysis
In this subsection, we investigate the impact of key parameters in the proposed heuristic methods, including (1) the threshold $\theta_{p}$ in NP, and (2) threshold $\xi$ in LP. All experiments are conducted on the network with 20 nodes, as described in Sec. VII-C2. GA is employed for task allocation, using the same configuration detailed in Sec. VII-C2.
@@ VII-D2 Threshold $\xi$
LP selects all nodes in the top $\xi+1$ levels as workers. In the special case when $\xi=0$, no nodes are selected, resulting in all computations being conducted locally at the master. This leads to the highest cost $J$, as shown in Fig. 11(a). As $\xi$ increases and more nodes are selected, performance improves, as indicated by the decreasing cost $J$. However, the performance improvement slows down when $\xi$ exceeds 3. The best performance is achieved when $\xi$ reaches its maximum value, $H$ (height of the network tree), which is 12 in this experiment. Given the rapid increase in execution time with higher $\xi$, as shown in Fig. 11(b), a proper value for $\xi$ can be chosen at the point where the rate of decrease in $J$ slows down.
@@ VIII Conclusion and Future Works
This paper introduces a novel multi-hop multi-layer distributed computing framework that expands the computing capabilities of networked computing systems. Unlike conventional distributed computing paradigms that limit resource sharing to one-hop neighborhoods, our framework explores layered network structures and multi-hop routing to enable resource sharing beyond one-hop neighborhoods, effectively utilizing the resources of the entire network. To optimize system performance, we formulated an MIP problem that jointly minimizes task completion time and energy consumption through optimal task allocation and scheduling. Two exact methods, CMO and PMO, were proposed to solve this problem optimally, with PMO enhancing CMO’s efficiency by exploiting the parallelism in task distribution across the master’s subtrees. We also introduced an offline-online computation scheme to enable the real-time execution of CMO and PMO and allow them to handle dynamic networks with time-varying characteristics. To further enhance efficiency and scalability, three heuristic methods were introduced, including NP and LP for reducing the number of workers and GA for efficiently finding (sub-)optimal solutions. Simulation results demonstrate the superiority of our approaches over existing distributed computing and computation offloading schemes. Moreover, PMO shows promising performance in both optimality and efficiency for small networks. For larger networks, the results suggest applying NP or LP to reduce workers before using GA or PMO for task allocation. Additionally, the results show that NP outperforms LP in terms of optimality but is less efficient due to its more rigorous worker selection process. Lastly, studies on the impact of NP’s and LP’s parameters offer insights into their configurations. In the future, we will extend this work to consider multi-task scenarios as well as dynamic and mobile networks. We will also investigate the hierarchical master-work paradigm.
Acknowledgement
We would like to thank National Science Foundation under Grant CAREER-2048266 for the support of this work.
@@ Acknowledgement
[Uncaptioned image] Ke Ma received the B.Sc. degree in Communication Engineer from Donghua University, Shanghai, China, in 2019 and the M.S. degree in Computer Engineer from University of California at Riverside, California, in 2022.

### v2

@@ VII Simulation Studies
In this section, we conduct simulation studies to evaluate the performance of the proposed approaches. We start by describing the experiment setup in Sec. VII-A. Next, we conduct two sets of studies to evaluate the optimality and efficiency of the proposed approaches in Sec. VII-B and Sec. VII-C, respectively. We then investigate the impact of key parameters in Sec. VII-D, followed by an analysis of the effects of network characteristics.
@@ VII-A Experiment Setup
We evaluate the proposed approaches on different network graphs generated using the method introduced in [erdds1959random]. The network graphs are then transformed into tree topologies by using Dijkstra’s algorithm. In each network topology, we configure the computing capacities $f_{i}$ of the servers by randomly generating values from the range of $[1,10]$GHz. The values of the data transmission rates $R_{ij}$ are randomly generated from the range of $[10,100]$Gbps. Moreover, we set $\gamma_{i}=10^{-2}$, and $e_{i}=30$dBm for all $i\in\mathcal{N}$. The task size is set to $Y=1$ Gbits and $b=10^{6}$ cycles/Gbit. All approaches are implemented in Python and evaluated on an Alienware Aurora R15 Gaming Desktop with a 12 Gen Intel i9 CPU and 64G of memory.
@@ VII-B Optimality Analysis
In the second experiment, we randomly set the weights to $w_{1}=0.5$ and $w_{2}=0.05$, so that both computation efficiency and energy consumption are considered in our approaches. Note that these weight values are also used in the following experiments. Fig. 5(b) shows the comparison results, demonstrating the promising performance of our approaches in balancing task completion time and energy consumption. In Fig. 6(a) and Fig. 6(b), we show the task completion time, i.e., $\max_{i\in\mathcal{N}}T_{i}^{total}$, and the maximum energy consumption by any server, i.e., $\max_{i\in\mathcal{N}}E_{i}^{total}$, respectively. The result indicates that our approach outperforms all the benchmarks in both task completion time and maximum energy consumption.
Figure 5: Total cost $J$ of different methods when considering (a) only time consumption; and (b) both time and energy consumption.
Figure 6: (a) Task completion time (b) Maximum energy consumption of different methods.
@@ VII-C1 Small-Scale Networks
The base-10 logarithm of the execution time, i.e., $\log_{10}T_{\text{exe}}$, of each method is shown in Fig. 7(b). As expected, the optimal methods, CMO and PMO, are more time-consuming than the three heuristic methods. Moreover, PMO, being a parallelized version of CMO, significantly reduces execution time in Topologies 2-4 due to its parallelism. For Topology 1, since the root has a single subtree, PMO is equivalent to CMO. Among the heuristic methods, LP achieves the least execution time by pruning the most nodes and significantly reducing the search space. GA, on the other hand, is the least efficient and even underperforms PMO in Topologies 2 & 4. This suggests that for small networks, PMO can be directly applied. Furthermore, comparing NP and PMO, we can observe that NP does not improve efficiency in all scenarios, despite reducing the number of workers. This is because only one node is “pruned” in each topology, and the time saved by pruning is offset by the overhead generated by the pruning procedure.
@@ VII-D Parameter Impact Analysis
In this subsection, we investigate the impact of key parameters in the proposed heuristic methods, including (1) the threshold $\theta_{p}$ in NP, (2) threshold $\xi$ in LP. All experiments are conducted on the network with 20 nodes, as described in Sec. VII-C2. GA is employed for task allocation, using the same configuration detailed in Sec. VII-C2.
@@ VII-D2 Threshold $\xi$
LP selects all nodes in the top $\xi$ levels as workers. In the special case when $\xi=0$, no nodes are selected, resulting in all computations being conducted locally at the master. This leads to the highest cost $J$, as shown in Fig. 11(a). As $\xi$ increases and more nodes are selected, performance improves, as indicated by the decreasing cost $J$. However, the performance improvement slows down when $\xi$ exceeds 3. The best performance is achieved when $\xi$ reaches its maximum value, $H$ (height of the network tree), which is 12 in this experiment. Given the rapid increase in execution time with higher $\xi$, as shown in Fig. 11(b), a proper value for $\xi$ can be chosen at the point where the rate of decrease in $J$ slows down.
@@
VII-E Impact of Network Characteristics
The optimal task distribution decisions highly rely on the network characteristics. In this subsection, we explore how communication and computing parameters, specifically $R_{i,j}$ and $f_{i}$, affect these decisions. For this analysis, we focus on Topology 4, as shown in 4 in Fig. 4.
In the first experiment, we vary $R_{0,1}$, which represents the communication capacity between the master node (Node 0) and its left child (Node 1), from 0.3 Gbps to 10 Gbps. All other settings remain consistent with the previous studies. The optimal task allocation derived by PMO is shown in Fig.12(a). As the figure demonstrates, when $R_{0,1}$ is small, communication becomes a bottleneck, preventing the master from offloading tasks to Node 1 or to its descendants. However, as $R_{0,1}$ increases, more workload is offloaded to Node 1. Once $R_{0,1}$ exceeds certain thresholds, tasks are also offloaded to Node 1’s children and even grandchildren. With more nodes contributing to workload distribution, nodes in the right subtree of the master begin to receive fewer tasks. This study suggests that if a communication link is too slow, both the connected downstream node and its descendants may be pruned from the topology before executing CMO/PMO. To understand the impact of the computing characteristic, we instead vary the computing power of Node 1, $f_{1}$, from 0.022 GHz to 21 GHz. As shown in Fig. 12(b), the master starts to offload tasks to Node 1 when its computing power exceeds a certain threshold. Moreover, when it shares more workload, the workloads assigned to all other nodes decrease simultaneously.
Notably, the network characteristics determine whether tasks are offloaded to a node, regardless of the total task size $Y$, as inferred from Theorem 2. To demonstrate this, we vary $Y$ while keeping the network characteristics constant. Table 1 summarizes the optimal task allocations and the corresponding total costs computed by PMO. As shown, when $Y$ increases, both the workload assigned to each node $y_{i}$ and the total cost $J$ rise proportionally.
Figure 12: Optimal task allocations at different values of (a) $R_{0,1}$ and (b) $f_{1}$.
Table I: Optimal task allocation and total cost for different values of $Y$
@@ VIII Conclusion and Future Works
This paper introduces a novel multi-layered distributed computing framework that expands the computing capabilities of networked computing systems. Unlike conventional distributed computing paradigms that limit resource sharing to one-hop neighborhoods, our framework explores layered network structures to enable resource sharing beyond one-hop neighborhoods, effectively utilizing the resources of the entire network. To optimize system performance, we formulated an MIP problem that jointly minimizes task completion time and energy consumption through optimal task allocation and scheduling. Two exact methods, CMO and PMO, were proposed to solve this problem optimally, with PMO enhancing CMO’s efficiency by exploiting the parallelism in task distribution across the master’s subtrees. We also introduced an offline-online computation scheme to enable the real-time execution of CMO and PMO and allow them to handle dynamic networks with time-varying characteristics. To further enhance efficiency and scalability, three heuristic methods were introduced, including NP and LP for reducing the number of workers and GA for efficiently finding (sub-)optimal solutions. Simulation results demonstrate the superiority of our approaches over existing distributed computing and computation offloading schemes. Moreover, PMO shows promising performance in both optimality and efficiency for small networks. For larger networks, the results suggest applying NP or LP to reduce workers before using GA or PMO for task allocation. The results also show that NP outperforms LP in terms of optimality but is less efficient due to its more rigorous worker selection process. Additionally, studies on the impact of NP’s and LP’s parameters offer insights into their configurations. Lastly, the analysis of network characteristics highlights how the communication and computing capacities of individual servers influence task distribution decisions. In the future, we will extend this work to consider multi-task scenarios as well as dynamic and mobile networks. We will also investigate the hierarchical master-work paradigm.
Acknowledgment
We would like to thank National Science Foundation under Grant CAREER-2048266 and CCF-2402689 for the support of this work.
@@ Acknowledgment
[Uncaptioned image] Ke Ma is currently a PhD candidate in the joint doctoral program of University of California, San Diego and San Diego State University. He received the B.Sc. degree in Communication Engineer from Donghua University, Shanghai, China, in 2019 and the M.S. degree in Computer Engineer from University of California at Riverside, California, in 2022.